Angle E and angle T and supplementary. If angle E is 23 degrees more than twice the complement of T, what are the two angles?
Answer by stanbon(75887)

Angle E and angle T and supplementary.
E = 180-T
---------------------
If angle E is 23 degrees more than twice the complement of T, what are the two angles?
E = 2(90-T) + 23
-------
Substitute for E and solve for T:
180-T = 2(90-T)+23
180 - T = 180 - 2T + 23
T = 23 degrees
---
E = 180-T = 157 degrees
=============================
Cheers,
Stan H.
=============================